### v2

# references.bib

Abstract
Image-event joint depth estimation methods leverage complementary modalities for robust perception, yet face challenges in generalizability stemming from two factors: 1) limited annotated image-event-depth datasets causing insufficient cross-modal supervision, and 2) inherent frequency mismatches between static images and dynamic event streams with distinct spatiotemporal patterns, leading to ineffective feature fusion. To address this dual challenge, we propose Frequency-decoupled Unified Self-supervised Encoder (FUSE) with two synergistic components: The Parameter-efficient Self-supervised Transfer (PST) establishes cross-modal knowledge transfer through latent space alignment with image foundation models, effectively mitigating data scarcity by enabling joint encoding without depth ground truth. Complementing this, we propose the Frequency-Decoupled Fusion module (FreDFuse) to explicitly decouple high-frequency edge features from low-frequency structural components, resolving modality-specific frequency mismatches through physics-aware fusion. This combined approach enables FUSE to construct a universal image-event encoder that only requires lightweight decoder adaptation for target datasets. Extensive experiments demonstrate state-of-the-art performance with $14\%$ and $24.9\%$ improvements in Abs.Rel on MVSEC and DENSE datasets. The framework exhibits remarkable zero-shot adaptability to challenging scenarios including extreme lighting and motion blur, significantly advancing real-world deployment capabilities. The source code for our method is publicly available at: https://github.com/sunpihai-up/FUSE.
I Introduction
Modern monocular depth estimation (MDE) increasingly utilizes multi-modal sensing to address the limitations of single-sensor systems [ramnet, xu2024unveiling], particularly by integrating conventional image sensors with emerging event cameras. Traditional image sensors, which operate using fixed-exposure light integration, effectively preserve structural information but suppress high-frequency temporal variations. This characteristic makes them suitable for static scene perception but limits their performance in dynamic scenarios. In contrast, event cameras asynchronously detect pixel-wise brightness changes with microsecond resolution [zhu2018ev], making them highly effective for capturing high-frequency motion but incapable of detecting static structures with subthreshold intensity gradients. The inherent complementarity in the frequency domain between these two modalities highlights the potential of image-event fusion as a promising approach for achieving robust depth estimation in dynamic environments.
Figure 1: Demonstration of FUSE’s superior performance in challenging conditions. (a) and (c) show the input image and event data, with blue/red in (c) indicating brightness decrease/increase. When image data fails due to low light and blur, event data provides complementary dynamic information. Our method (d) leverages multimodal synergy, recovering the traffic light missed by the state-of-the-art image depth model [depthanything2], as highlighted in the orange box.
Figure 2: Overview of the FUSE framework. FUSE integrates an image encoder, event encoder, Frequency-Decoupled Fusion module (FreDFuse), and depth decoder. The image encoder and depth decoder are initialized with a pre-trained MDE model. A two-stage knowledge transfer strategy fine-tunes the event encoder and FreDFuse. In Stage I, the event encoder is initialized with the image encoder’s weights, and only the LoRA matrix and Patch Embed are fine-tuned using clean event data. In Stage 2, randomly degraded image-event pairs are used, and only the FreDFuse is fine-tuned. The MDE model supervises both stages with clean image data in the output and latent spaces.
Recent studies [ramnet, srfnet, pcdepth] have explored leveraging the complementary characteristics of image and event modalities for joint MDE. These methods typically encode both modalities into a unified feature space, followed by customized feature fusion modules for depth prediction. However, their reliance on supervised learning with depth ground truth from image-event pairs has revealed poor generalization capabilities. This limitation arises from the scarcity and limited diversity of existing labeled image-event datasets [mvsec] compared to large-scale image-depth datasets such as NYUv2 [nyuv2] and KITTI [kitti].
Additionally, although the frequency-domain complementarity between images and events presents theoretical advantages, mismatches in their frequency characteristics introduce significant challenges for feature fusion. Conventional fusion strategies risk destructive interference [chen2024frequency]: high-frequency event features can disrupt the structural continuity of image representations, while low-frequency image components may suppress critical motion cues captured by events. This inherent frequency divergence, combined with limited training data, poses a dual challenge—requiring models to bridge the modality gap with insufficient supervision and resolve conflicting frequency signatures during feature integration.
Inspired by the impressive generalization capabilities of image-based depth foundation models (e.g., Depth Anything [depthanything1]), which are trained on millions of images, we propose a potential solution: transferring their well-learned image-depth priors to bridge the gap between event-based and depth-based representations. This insight motivates our two-fold innovation: 1) addressing data scarcity through parameter-efficient transfer learning, and 2) resolving intrinsic frequency mismatches via frequency decoupling mechanisms.
We introduce FUSE, a frequency-decoupled, unified, self-supervised encoder-based method designed to transfer knowledge from image-based depth models to image-event joint (MDE) without requiring depth ground truth. FUSE consists of two main components: 1) Parameter-Efficient Self-Supervised Transfer (PST): A two-phase training strategy with decoupled feature alignment and fusion stages. In Phase I, an event encoder is trained using lightweight adapter tuning to establish a shared latent space between image and event modalities. Phase II focuses on frequency-aware fusion training, incorporating a novel fusion module, FreDFuse, under simulated sensor degradation conditions; 2) Frequency-Decoupled Fusion Module (FreDFuse): Serving as the core of PST’s second phase, this module separates features into distinct frequency components before fusion. By utilizing frequency-specific processing pipelines, it enables complementary integration while minimizing inter-frequency interference, thus overcoming the limitations of traditional single-bandwidth fusion approaches.
To validate the effectiveness of FUSE, we conduct extensive experiments on the MVSEC and DENSE datasets, demonstrating state-of-the-art (SOTA) performance in MDE. As shown in Fig. 1, FUSE outperforms the current SOTA image depth model [depthanything2] under zero-shot settings, particularly in night-time driving scenarios affected by motion blur, highlighting its robustness. In summary, our contributions are as follows:
  • We propose FUSE, the first self-supervised framework for generalizable image-event depth estimation that transfers knowledge from image-only depth foundation models without requiring depth-labeled image-event pairs. This addresses the critical challenge of scarce annotated image-event-depth datasets.
  • Parameter-Efficient Self-Supervised Transfer (PST): A novel two-stage training strategy that bridges the image-event representation gap through lightweight LoRA-based tuning and degradation-robust distillation, significantly reducing trainable parameters compared to full fine-tuning.
  • Frequency-Decoupled Fusion Module (FreDFuse): A frequency-aware fusion module that explicitly disentangles high-frequency edge dynamics (from events) and low-frequency structural priors (from images). This approach resolves modality-specific frequency mismatches while preserving complementary cues.
II Related work
II-A Image-based Monocular Depth Estimation
Deep learning-based methods have become the dominant paradigm in MDE. Eigen et al. [eigen] were the first to introduce a multi-scale fusion network for depth prediction through regression. Subsequent studies have improved depth estimation accuracy by reformulating the regression task as classification [adabins, binsformer, shao2023iebins], introducing novel network architectures [dpt], and incorporating additional prior knowledge [nddepth, newcrf]. Despite these advancements, the generalization capability of such models remains a major challenge.To address this limitation, recent research [midas, metric3d] has focused on training models on large-scale datasets aggregated from diverse sources, leading to substantial improvements. A milestone in this direction is MiDaS [midas], which mitigates the impact of varying depth scales across datasets by employing an affine-invariant loss. More recently, the DepthAnything series [depthanything1, depthanything2] has further enhanced the utilization of unlabeled data alongside large-scale labeled datasets, achieving remarkable performance. However, due to the inherent limitations of traditional image cameras, image-based MDE methods are still susceptible to factors such as lighting conditions and motion blur, which can degrade their performance [imgbad1, imgbad2, imgbad3].
II-B Event-based Monocular Depth Estimation
Event-based MDE methods have attracted attention for their effectiveness under extreme lighting and high-speed motion conditions, making them suitable for applications in autonomous driving and robotic navigation [eventsurvey]. E2Depth [e2depth] was the first method to propose dense MDE using event data, employing a recurrent encoder-decoder architecture. Mixed-EF2DNet [mixed] introduced a flow network to capture temporal information effectively, while EReFormer [ereformer] leveraged a Transformer-based architecture for improved performance. Additionally, HMNet [hmnet] incorporated multi-level memory units to handle long-term dependencies efficiently. Nonetheless, these methods continue to face challenges due to the sparsity of event data and the limited availability of training datasets [eventbad].
II-C Image-Event Fusion
The complementary characteristics of event and image data have been leveraged to enhance various visual tasks, including semantic segmentation [zhang2023cmx, zhang2023cmx], object detection [zhang2021object, zhou2023rgb], and depth estimation [ramnet, srfnet]. RAMNet [ramnet] employs an RNN-based approach to effectively utilize asynchronous event data, while EVEN [shi2023even] improves low-light performance by fusing modalities prior to inputting them into an MDE network. SRFNet [srfnet] guides modality interaction using spatial reliability masks, and PCDepth [pcdepth] integrates features at the modality level for more effective fusion. Our FUSE is the first to leverage knowledge transfer from an image foundation model for both feature encoding and fusion, significantly enhancing robustness and accuracy in image-event joint depth estimation.
III Methodology
III-A Overview
In this section, we introduce FUSE, a low-cost and robust transfer learning framework for generalizable depth prediction. An overview of the FUSE architecture is provided in Fig. 2, our framework consists of four key components: an event encoder $\mathcal{E}_{E}$, an image encoder $\mathcal{E}_{I}$, a Frequency-Decoupled Fusion module (FreDFuse), and a depth decoder $\mathcal{D}$. Given an input image $\mathbf{I}$ and its event stream $\mathbf{S}$, the target is to predict the dense depth map $\mathbf{d}^{*}$. Our FUSE provides a powerful image-event joint encoder ($\mathcal{E}_{I}$, $\mathcal{E}_{E}$, FreDFuse) with exceptional generalization capabilities, enabling robust performance with minimal adaptation of the depth decoder for target datasets.
In Sec.III-B, we first describe the event processing pipeline. Subsequently, in Sec.III-C, we detail the process of obtaining the image-event joint encoder via PST. Finally, in Sec. III-D, we present the FreDFuse in detail.
Figure 3: Overview of our FreDFuse. FreDFuse decouples image features $\mathbf{F}_{I}$ and event features $\mathbf{F}_{E}$ into high- and low-frequency components using a Gaussian-Laplacian pyramid. Multi-scale features are fused top-down with $1\times 1$ group convolutions and channel shuffle. Fusion in the high-frequency branch is event-driven, while the low-frequency branch is image-driven. The high- and low-frequency components are combined through addition and processed by LayerNorm.
TABLE I: Quantitative results on the MVSEC datast [mvsec]. $\downarrow$ indicates lower is better and denotes higher is better. The best results are highlighted in bold, while the second-best outcomes are underlined.
[…]
III-B Event Representation
The event stream is first transformed into a voxel grid representation [ramnet] to make it compatible with existing network architectures. Each event $e_{n}$ is represented by the tuple $(x_{n},y_{n},t_{n},p_{n})$, which means that at time $t_{n}$, the log intensity change at the pixel position $(x_{n},y_{n})$ exceeds a threshold. The polarity $p_{n}\ (\pm 1)$ indicates whether the intensity increases or decreases. As such, the asynchronous event stream $\mathbf{S}=\{e_{n}\}_{n=0}^{N-1}$ must be transformed into a frame-like representation to be compatible with existing network architectures. Following prior works [ramnet], [srfnet], [pcdepth], [e2depth] we convert the temporal stream of events into a voxel grid [gehrig2019end].
Given an event stream $\mathit{S}=\{e_{n}\}_{n=0}^{N-1}$ within the time interval $\Delta T=t_{N-1}-t_{0}$, we transform it into a voxel grid with spatial dimensions $H\times W$ and $B$ time bins:
$\mathit{V}(x,y,t)=\sum p_{i}\delta\left(x-x_{i},y-y_{i}\right)\max\left\{0,1-\left|t-t_{i}^{*}\right|\right\}$ (1)
where $t_{i}^{*}=\frac{B-1}{\Delta T}(t_{i}-t_{0})$. The voxel grid $\mathit{V}\in\mathbb{R}^{H\times W\times B}$. We choose $B=3$ time bins.
The similar structural form facilitates knowledge transfer from image-based models, enabling the leveraging of pre-trained networks for enhanced performance. At the same time, the image data can be used without additional processing.
III-C Parameter-efficient Self-supervised Transfer
Our Parameter-efficient Self-supervised Transfer (PST) strategy addresses the critical challenge of event-based depth estimation data scarcity by establishing a unified foundation model for image-event joint MDE through efficient cross-modal knowledge transfer. This self-supervised approach eliminates dependency on expensive depth ground truth while achieving two key objectives: 1) transferring image domain priors to the image-event joint domain using only paired image-event data, and 2) preserving the foundation model’s generalization capability through parameter-efficient adaptation. As shown in Fig. 2, PST operates through two cascaded stages:
Stage I: Parameter-Efficient Feature Alignment: In this stage, we align event data representations with the image domain’s latent space. We initialize $\mathcal{E}_{E}$ and $\mathcal{D}_{E}$ using weights from Depth Anything V2 [depthanything2] to inherit its geometric understanding ability. To enable efficient adaptation while preventing catastrophic forgetting [forgetcatastrophic], we implement LoRA[lora], where only the LoRA matrices and PatchEmbed Layer parameters (1.8% of $\mathcal{E}_{E}$) are updated. This allows adaptation to event data while retaining the foundation model’s generalization capability. The alignment is driven by a compound objective function (detailed in Sec. III-E) that enforces consistency in both output space (depth prediction) and latent representations.
Stage II: Robust Feature Fusion: In this stage, we focus on training the FreDFuse (Sec. III-D) in the image-event joint estimator to robustly fuse multi-modal features from image and event. The image encoder $\mathcal{E}_{I}$ and decoder $\mathcal{D}$ are inherited from the image foundation model, e.g., Depth Anything V2, while the event encoder $\mathcal{E}_{E}$ is obtained from Stage I. Only the FreDFuse is trainable during this stage. To improve the robustness and generalization of the proposed method, we also generate randomly degraded image-event pairs as inputs during training. The degraded inputs incorporate diverse noise types, including brightness adjustment, overexposure, motion blur, and occlusion masking. This systematic degradation strategy compels the feature fusion module to identify robust cross-modal relationships that withstand various sensor-specific degradations, thereby enhancing the model’s generalizability and resilience. Here, we use the same objective as Eq. 10 to maintain cross-modal consistency.
III-D Frequency-Decoupled Fusion Module
To leverage the complementary characteristics of different visual sensors, we propose a frequency-decoupled fusion method. Image sensors excel at capturing low-frequency static scenes, while event cameras are adept at detecting high-frequency dynamic changes. Our Frequency-Decoupled Feature Integration module (FreDFuse) leverages these complementary strengths by processing and fusing features in separate frequency domains, enabling more effective multi-modal depth estimation. Fig. 3 shows the overview of FreDFuse.
Frequency Decoupling via Gaussian-Laplacian Pyramids: Given token sequences from image features $\mathbf{F}_{I}\in\mathbb{R}^{B\times N\times C}$ and event features $\mathbf{F}_{E}\in\mathbb{R}^{B\times N\times C}$, we first decompose them into low-frequency and high-frequency components using Gaussian-Laplacian pyramids. For each modality $m\in\{I,E\}$, the decomposition is formulated as:
$\mathbf{G}_{m},\mathbf{L}_{m}=\mathcal{P}_{m}(\mathbf{F}_{m}),$ (2)
where $\mathcal{P}_{m}(\cdot)$ denotes the Gaussian-Laplacian decoupling operator. Specifically, the input tokens are reshaped into 2D feature maps and processed through multi-scale pyramids. The Gaussian pyramid $\mathbf{G}_{m}$ captures low-frequency components via iterative blurring and downsampling:
$\mathbf{G}_{m}^{(l)}=\text{Down}(\text{GaussianBlur}(\mathbf{G}_{m}^{(l-1)})),$ (3)
where $l\in\{1,\dots,L\},L=3$ denotes the pyramid level. The Laplacian pyramid $\mathbf{L}_{m}$, encoding high-frequency details, is derived by subtracting upsampled Gaussian levels:
$\mathbf{L}_{m}^{(l)}=\mathbf{G}_{m}^{(l)}-\text{Up}(\mathbf{G}_{m}^{(l+1)}).$ (4)
Top-Down Fusion of Multi-Level Pyramid Features: For each pyramid ($\mathbf{G}_{I},\mathbf{L}_{I},\mathbf{G}_{E},\mathbf{L}_{E}$), multi-scale features are fused by upsampling and concatenating layers, followed by grouped convolutions:
$\mathbf{P}_{\text{cat}}^{(l)}=\text{Concat}(\mathbf{P}^{(l-1)},\text{Up}(\mathbf{P}^{(l)})),$ (5)
$\mathbf{P}_{\text{comp}}^{(l)}=\text{GroupConv}_{1\times 1}(\mathbf{P}_{\text{cat}}^{(l)}),$ (6)
$\mathbf{P}_{\text{fused}}^{(l)}=\text{GroupConv}_{1\times 1}(\text{ChannelShuffle}(\mathbf{F}_{\text{comp}}^{(l)})).$ (7)
Cross-Branch Frequency-Guided Fusion: The decoupled features are fused in two branches, each guided by the dominant modality in the target frequency band. For the high-frequency branch, $\mathbf{F}_{E}^{\text{high}}$ is the Query, while $\mathbf{F}_{I}^{\text{high}}$ is the Key and Value. For the low-frequency branch, $\mathbf{F}_{I}^{\text{low}}$ is the Query, while $\mathbf{F}_{E}^{\text{low}}$ is the Key and Value. Both branches use multi-head cross-attention:
$\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}}\right)\mathbf{V},$ (8)
where $d$ is the dimension of each attention head. The fused features are recombined and projected to the original space:
$\mathbf{F}_{\text{fused}}=\text{LayerNorm}\left(\text{Linear}(\mathbf{F}_{\text{low}}+\mathbf{F}_{\text{high}})\right).$ (9)
III-E Optimization Objectives
Optimization Objective of PST: The PST framework adopts a compound loss to achieve two critical objectives: 1) prediction-level distillation via output-space alignment, and 2) latent-space regularization to preserve transferable representations. Given clean image $\mathbf{I}$, the image foundation model generates pseudo depth $\mathbf{d}$ and latent features $\mathbf{F}$, which provide supervision for Stage I and Stage II. The depth map and latent features of the output of Stage I and Stage II are defined as $\mathbf{d^{*}}$ and $\mathbf{F}^{*}$. We formulate a compound loss function ${L}_{\text{align}}$:
${L}_{\text{align}}={L}_{1}+{L}_{\cos},$ (10)
where ${L}_{1}=||{d}-{d}^{*}||_{1}$ aligns output predictions, and ${L}_{\cos}=\left[1-\cos(\mathbf{F},\mathbf{F}^{*})\right]\cdot\mathbb{I}({\alpha\leq\cos(\mathbf{F},\mathbf{F}^{*})\leq\beta})$ aligns latent representations. Here, $\alpha=0.2$, $\beta=0.85$, and $\mathbb{I}(\cdot)$ ensure loss activation within the specified similarity range, preventing over-regularization.
Optimization Objective of Target Datasets: When transferring to target datasets with metric depth labels, we replace pseudo-supervision with the Scale-invariant Logarithmic (SiLog) loss [eigen]. Given a predicted depth map $\mathbf{d^{*}}$ and depth ground truth $\mathbf{d}$, the loss is computed over the pixels as follows:
${L}_{\text{SiLog}}=\sqrt{\frac{1}{N}\sum_{i}e_{i}^{2}-\lambda\left(\frac{1}{N}\sum_{i}e_{i}\right)^{2}},$ (11)
where $e_{i}=\ln d_{i}-\ln d^{*}_{i}$, $N$ is the total number of valid pixels.
Figure 4: Qualitative analysis of the MVSEC dataset, outdoor_day1 scene. (a) and (b) show the input image and event data; (c) depicts the image-event joint estimation by HMNet[hmnet]; (d) shows the event-only estimation by HMNet; (e) presents our proposed FUSE; (f) shows the depth ground truth.
Figure 5: Qualitative analysis of the MVSEC dataset, outdoor_night1 scene. (a) and (b) show the input image and event data; (c) depicts the image-event joint estimation by HMNet[hmnet]; (d) shows the event-only estimation by HMNet; (e) presents our proposed FUSE; (f) shows the depth ground truth.
Figure 6: Qualitative comparison of zero-shot predictions when the image modality is severely degraded.(a) denotes the original image, (b) represents the image degraded by low light, overexposure, and occlusion masks, (c) denotes the event data, and (d) and (e) show the predicted results from DepthAnything V2 [depthanything2] and our FUSE, respectively.
IV Experiment
To validate the effectiveness of our proposed FUSE framework in addressing the scarcity of event-depth data and modality disparity, we conduct comprehensive experiments under strictly controlled conditions. Our evaluation protocol emphasizes three critical aspects: (1) Generalization capability across real-world (MVSEC) and synthetic (DENSE) scenarios; (2) Performance constraints with frozen image-event joint encoders; (3) Computational efficiency in multi-modal fusion.
IV-A Experimental Settings
We utilize EventScape [ramnet] for image-event pairs in PST (without depth ground truth). We evaluate on MVSEC [mvsec] and DENSE [e2depth] datasets, where MVSEC is a real-world scenario while EventScape and DENSE are synthetic scenarios.
MVSEC Dataset: The MVSEC dataset uses a pair of DAVIS cameras to capture grayscale images and event data at a resolution of $346\times 260$. Depth ground truth is recorded using a LiDAR sensor at 20 Hz. Grayscale images are captured at 10 Hz during the day and 45 Hz at night.
Metrics: Following [pcdepth], we evaluate performance using the metrics: Absolute Relative Error (Abs.Rel), Root Mean Square Error (RMSE), Logarithmic Squared Error (RMSELog), average absolute depth errors at different cut-off depth distances (i.e., 10m, 20m, and 30m), and accuracy ($\delta<1.25^{n},n=1,2,3$).
Implementation Details: The model is trained using the Adam optimizer with a learning rate of 5e-5, implemented in PyTorch on two RTX 3090ti GPUs with a batch size of 48. PST on EventScape [ramnet] is trained for 10 epochs, and fine-tuning on MVSEC and DENSE for 20 epochs. Manual alignment is needed for the MVSEC dataset due to the asynchronous nature of images, events, and depth labels. As in [pcdepth], event data from the first 50 ms prior to each depth ground truth is paired with the most recent image to form the data pairs.
IV-B Evaluation on MVSEC and DENSE Datasets
We report quantitative results on the real-world MVSEC [mvsec] and synthetic DENSE [e2depth] datasets. For MVSEC, we use the outdoor_day2, outdoor_night2, and outdoor_night3 sequences for training, and outdoor_day1 and outdoor_night1 for testing, with depth range from 1.97 to 80 meters [pcdepth]. For DENSE, we use the provided training and test sets, with depth range from 3.34 to 1000 meters [ramnet, srfnet]. We freeze the image-event joint encoder and train only the depth decoder.
Analysis on MVSEC Dataset: Tab. I presents a quantitative comparison between our method and SOTA methods for event-based estimation and image-event joint estimation on the MVSEC dataset. In the outdoor_day1 scene, our method outperforms the best image-event joint estimation approach [pcdepth] by $14\%$ and $10.2\%$ in Abs.Rel and RMSELog, respectively. The progress achieved by freezing the encoder and dealing with the significant disparity between source and target data clearly demonstrates the effectiveness of our proposed FUSE in addressing the scarcity of event depth data and providing robust and accurate depth estimation. Fig. 4 and Fig. 5 present qualitative comparisons with HMNet[hmnet] for the outdoor_day1 and outdoor_night1 scenes, respectively. FUSE provides finer and more stable depth predictions. Compared to HMNet, our method demonstrates a superior ability to preserve structural details, delivering more consistent and accurate predictions on both buildings and vegetation.
Analysis on DENSE Dataset: Tab. II presents the quantitative results on the synthetic DENSE dataset [e2depth]. Our method achieves improvements of $24.9\%$ in Abs.Rel and $33.4\%$ in RMSELog over previous approaches. For the average absolute depth error at the truncated distance, our method ranks either the best or the second-best among the evaluated methods. Fig. 6 shows a qualitative comparison under extreme damage to the image modality. Our FUSE is still able to provide stable depth estimation even when some regions of the image are completely destroyed.
TABLE II: Comparison of performances on synthetic DENSE dataset [e2depth]
[…]
IV-C Ablation Studies
TABLE III: Ablation studies for Image-Event fusion models
  Knowledge Transformer Feature Fusion Module Trainable Paras (M) Baseline-1 None Cross-Attention 49.2 Baseline-2 None FreDFuse 51.6 Baseline-3 One-stage FreDFuse 6+2.7 FUSE PST FreDFuse 1.4+4.7+2.7
We perform ablation experiments using the ViT-Small model as the backbone on the MVSEC dataset, as outlined in Tab.III. In Baseline-1 and Baseline-2, the model is trained from scratch on image-event pairs with depth ground truth. Baseline-1 uses two cross-attention mechanisms for feature fusion, while Baseline-2 incorporates the proposed FreDFuse (Sec. III-D). Baseline-3 combines FreDFuse with a one-stage knowledge transfer approach, initializing the event encoder with image-based foundation model weights and training both the event encoder and FreDFuse simultaneously.
Tab. IV and Tab. V present the results of the ablation experiments on the MVSEC dataset, specifically in the outdoor_day1 and outdoor_night1 scenes. The results demonstrate that both components of our proposed method, PST and FreDFuse, positively influence model performance.
Effectiveness of FreDFuse: In the outdoor_day1 scene, Baseline-2 demonstrates an $8\%$ improvement over Baseline-1 in the Abs. Rel metric. Similar improvements are observed across other scenes and metrics. This performance gain can be attributed to FreDFuse’s ability to decouple features into high-frequency and low-frequency branches. By allowing the modality that excels in each branch to dominate the feature fusion, FreDFuse effectively mitigates destructive interference caused by frequency mismatches and enhances inter-modal complementation.
Effectiveness of PST: Compared to Baseline-2, FUSE integrates our proposed PST, which reduces the training parameters by $82.2\%$ while achieving an average performance improvement of $19.7\%$ across all metrics and scenarios. This finding suggests that PST can effectively leverage the knowledge from image-based foundation models to mitigate the data scarcity issue in image-event joint depth estimation. In comparison to Baseline-3, FUSE, which employs the complete two-stage knowledge transfer process, also shows performance improvements on most metrics in various scenarios. This highlights the significance of the two-stage training with degraded image-event pairs for enhancing the model’s generalization and robustness.
V Conclusion
In this paper, we propose FUSE, a novel framework that addresses two fundamental challenges in image-event joint depth estimation: cross-modal knowledge transfer under extreme data scarcity and physics-driven frequency conflicts during feature fusion. By aligning image-event latent spaces through a two-stage adapter tuning strategy, PST successfully transfers geometric priors from image depth foundation models to the joint modality domain without requiring any depth-labeled image-event pairs. This proves that event-depth representations can be effectively bootstrapped from image counterparts when guided by physics-aware constraints, opening new possibilities for data-efficient cross-modal learning. Unlike conventional fusion heuristics, the proposed FreDFuse explicitly resolves the spectral incompatibility between static images (low-frequency dominance) and dynamic events (high-frequency bias) through Gaussian-Laplacian pyramid decomposition. The FUSE framework achieves SOTA performance on both real-world (MVSEC) and synthetic (DENSE) benchmarks while enabling lightweight deployment.
While FUSE demonstrates strong zero-shot generalization, two limitations warrant further investigation. Current fusion treats events as fixed-interval histograms, potentially underutilizing their microsecond temporal resolution. The framework assumes synchronized image-event pairs, which may not hold for all hardware setups. Future work will explore event-guided dynamic fusion that adaptively weights frequency branches based on motion intensity, and unsupervised cross-sensor calibration to relax synchronization requirements. These extensions could further solidify image-event depth estimation as a universal solution for extreme environment perception.
TABLE IV: Quantitative results of the ablation experiments on the MVSEC[mvsec] outdoor_day1
[…]
10m 20m 30m
Baseline-1 0.612 0.366 8.451 0.453 1.711 2.754 3.303
Baseline-2 0.633 0.336 8.403 0.433 1.488 2.519 3.137
Baseline-3 0.704 0.233 6.690 0.313 1.076 1.802 2.210
FUSE 0.719 0.229 6.169 0.294 1.035 1.848 2.267
TABLE V: Quantitative results of the ablation experiments on the MVSEC[mvsec] outdoor_night1
[…]
10m 20m 30m
Baseline-1 0.525 0.356 7.698 0.413 1.899 2.634 3.077
Baseline-2 0.531 0.351 7.583 0.418 1.817 2.545 3.018
Baseline-3 0.611 0.268 6.873 0.354 1.326 2.025 2.543
FUSE 0.613 0.267 6.785 0.352 1.305 1.998 2.513
[…]

### v3

@@
-references.bib
+FUSE: Label-Free Image-Event Joint Monocular Depth Estimation via Frequency-Decoupled Alignment and Degradation-Robust Fusion
 Abstract
 Image-event joint depth estimation methods leverage complementary modalities for robust perception, yet face challenges in generalizability stemming from two factors: 1) limited annotated image-event-depth datasets causing insufficient cross-modal supervision, and 2) inherent frequency mismatches between static images and dynamic event streams with distinct spatiotemporal patterns, leading to ineffective feature fusion.
-To address this dual challenge, we propose Frequency-decoupled Unified Self-supervised Encoder (FUSE) with two synergistic components: The Parameter-efficient Self-supervised Transfer (PST) establishes cross-modal knowledge transfer through latent space alignment with image foundation models, effectively mitigating data scarcity by enabling joint encoding without depth ground truth.
-Complementing this, we propose the Frequency-Decoupled Fusion module (FreDFuse) to explicitly decouple high-frequency edge features from low-frequency structural components, resolving modality-specific frequency mismatches through physics-aware fusion.
+To address this dual challenge, we propose Frequency-decoupled Unified Self-supervised Encoder (FUSE) with two synergistic components:
+The Parameter-efficient Self-supervised Transfer (PST) leverages image foundation models for cross-modal knowledge transfer, effectively mitigating data scarcity by enabling joint encoding without depth ground truth.
+Complementing this, the Frequency-Decoupled Fusion module (FreDFuse) resolves modality-specific frequency mismatches by decoupling features into high- and low-frequency bands and then performing a guided cross-attention fusion, where the modality dominant in each band steers the integration.
 This combined approach enables FUSE to construct a universal image-event encoder that only requires lightweight decoder adaptation for target datasets.
 Extensive experiments demonstrate state-of-the-art performance with $14\%$ and $24.9\%$ improvements in Abs.Rel on MVSEC and DENSE datasets.
-The framework exhibits remarkable zero-shot adaptability to challenging scenarios including extreme lighting and motion blur, significantly advancing real-world deployment capabilities.
+The framework exhibits remarkable robustness and generalization in challenging scenarios, including extreme lighting and motion blur, significantly advancing its real-world deployment capabilities.
 The source code for our method is publicly available at: https://github.com/sunpihai-up/FUSE.
 I Introduction
-Modern monocular depth estimation (MDE) increasingly utilizes multi-modal sensing to address the limitations of single-sensor systems [ramnet, xu2024unveiling], particularly by integrating conventional image sensors with emerging event cameras.
+Modern monocular depth estimation (MDE) increasingly utilizes multi-modal sensing to address the limitations of single-sensor systems [6, 24], particularly by integrating conventional image sensors with emerging event cameras.
 Traditional image sensors, which operate using fixed-exposure light integration, effectively preserve structural information but suppress high-frequency temporal variations.
 This characteristic makes them suitable for static scene perception but limits their performance in dynamic scenarios.
-In contrast, event cameras asynchronously detect pixel-wise brightness changes with microsecond resolution [zhu2018ev], making them highly effective for capturing high-frequency motion but incapable of detecting static structures with subthreshold intensity gradients.
+In contrast, event cameras asynchronously detect pixel-wise brightness changes with microsecond resolution [32], making them highly effective for capturing high-frequency motion but incapable of detecting static structures with subthreshold intensity gradients.
 The inherent complementarity in the frequency domain between these two modalities highlights the potential of image-event fusion as a promising approach for achieving robust depth estimation in dynamic environments.
@@
 When image data fails due to low light and blur, event data provides complementary dynamic information.
-Our method (d) leverages multimodal synergy, recovering the traffic light missed by the state-of-the-art image depth model [depthanything2], as highlighted in the orange box.
+Our method (d) leverages multimodal synergy, recovering the traffic light missed by the state-of-the-art image depth model [25], as highlighted in the orange box.
 Figure 2: Overview of the FUSE framework.
@@
 The MDE model supervises both stages with clean image data in the output and latent spaces.
-Recent studies [ramnet, srfnet, pcdepth] have explored leveraging the complementary characteristics of image and event modalities for joint MDE.
+Recent studies [6, 16, 13] have explored leveraging the complementary characteristics of image and event modalities for joint MDE.
 These methods typically encode both modalities into a unified feature space, followed by customized feature fusion modules for depth prediction.
 However, their reliance on supervised learning with depth ground truth from image-event pairs has revealed poor generalization capabilities.
-This limitation arises from the scarcity and limited diversity of existing labeled image-event datasets [mvsec] compared to large-scale image-depth datasets such as NYUv2 [nyuv2] and KITTI [kitti].
+This limitation arises from the scarcity and limited diversity of existing labeled image-event datasets [33] compared to large-scale image-depth datasets such as NYUv2 [23] and KITTI [8].
 Additionally, although the frequency-domain complementarity between images and events presents theoretical advantages, mismatches in their frequency characteristics introduce significant challenges for feature fusion.
-Conventional fusion strategies risk destructive interference [chen2024frequency]: high-frequency event features can disrupt the structural continuity of image representations, while low-frequency image components may suppress critical motion cues captured by events.
+Conventional fusion strategies risk destructive interference [1]: high-frequency event features can disrupt the structural continuity of image representations, while low-frequency image components may suppress critical motion cues captured by events.
 This inherent frequency divergence, combined with limited training data, poses a dual challenge—requiring models to bridge the modality gap with insufficient supervision and resolve conflicting frequency signatures during feature integration.
-Inspired by the impressive generalization capabilities of image-based depth foundation models (e.g., Depth Anything [depthanything1]), which are trained on millions of images, we propose a potential solution: transferring their well-learned image-depth priors to bridge the gap between event-based and depth-based representations.
+Inspired by the impressive generalization capabilities of image-based depth foundation models (e.g., Depth Anything [26]), which are trained on millions of images, we propose a potential solution: transferring their well-learned image-depth priors to bridge the gap between event-based and depth-based representations.
 This insight motivates our two-fold innovation: 1) addressing data scarcity through parameter-efficient transfer learning, and 2) resolving intrinsic frequency mismatches via frequency decoupling mechanisms.
@@
 As shown in Fig.
-1, FUSE outperforms the current SOTA image depth model [depthanything2] under zero-shot settings, particularly in night-time driving scenarios affected by motion blur, highlighting its robustness.
+1, FUSE outperforms the current SOTA image depth model [25] under zero-shot settings, particularly in night-time driving scenarios affected by motion blur, highlighting its robustness.
 In summary, our contributions are as follows:
@@
 Eigen et al.
-[eigen] were the first to introduce a multi-scale fusion network for depth prediction through regression.
-Subsequent studies have improved depth estimation accuracy by reformulating the regression task as classification [adabins, binsformer, shao2023iebins], introducing novel network architectures [dpt], and incorporating additional prior knowledge [nddepth, newcrf].
-Despite these advancements, the generalization capability of such models remains a major challenge.To address this limitation, recent research [midas, metric3d] has focused on training models on large-scale datasets aggregated from diverse sources, leading to substantial improvements.
-A milestone in this direction is MiDaS [midas], which mitigates the impact of varying depth scales across datasets by employing an affine-invariant loss.
-More recently, the DepthAnything series [depthanything1, depthanything2] has further enhanced the utilization of unlabeled data alongside large-scale labeled datasets, achieving remarkable performance.
-However, due to the inherent limitations of traditional image cameras, image-based MDE methods are still susceptible to factors such as lighting conditions and motion blur, which can degrade their performance [imgbad1, imgbad2, imgbad3].
+[3] were the first to introduce a multi-scale fusion network for depth prediction through regression.
+Subsequent studies have improved depth estimation accuracy by reformulating the regression task as classification [4, 12], introducing novel network architectures [17], and incorporating additional prior knowledge [20, 29].
+Despite these advancements, the generalization capability of such models remains a major challenge.To address this limitation, recent research [18] has focused on training models on large-scale datasets aggregated from diverse sources, leading to substantial improvements.
+A milestone in this direction is MiDaS [18], which mitigates the impact of varying depth scales across datasets by employing an affine-invariant loss.
+More recently, the DepthAnything series [26, 25] has further enhanced the utilization of unlabeled data alongside large-scale labeled datasets, achieving remarkable performance.
+Concurrently, other research [27] has focused on improving generalization through model-centric approaches, such as employing stabilized adversarial training to boost robustness against domain shifts.
+However, due to the inherent limitations of traditional image cameras, image-based MDE methods are still susceptible to factors such as lighting conditions and motion blur, which can degrade their performance [2].
 II-B Event-based Monocular Depth Estimation
-Event-based MDE methods have attracted attention for their effectiveness under extreme lighting and high-speed motion conditions, making them suitable for applications in autonomous driving and robotic navigation [eventsurvey].
-E2Depth [e2depth] was the first method to propose dense MDE using event data, employing a recurrent encoder-decoder architecture.
-Mixed-EF2DNet [mixed] introduced a flow network to capture temporal information effectively, while EReFormer [ereformer] leveraged a Transformer-based architecture for improved performance.
-Additionally, HMNet [hmnet] incorporated multi-level memory units to handle long-term dependencies efficiently.
-Nonetheless, these methods continue to face challenges due to the sparsity of event data and the limited availability of training datasets [eventbad].
+Event-based MDE methods have attracted attention for their effectiveness under extreme lighting and high-speed motion conditions, making them suitable for applications in autonomous driving and robotic navigation [5].
+E2Depth [10] was the first method to propose dense MDE using event data, employing a recurrent encoder-decoder architecture.
+Mixed-EF2DNet [21] introduced a flow network to capture temporal information effectively, while EReFormer [14] leveraged a Transformer-based architecture for improved performance.
+Additionally, HMNet [9] incorporated multi-level memory units to handle long-term dependencies efficiently.
+Nonetheless, these methods continue to face challenges due to the sparsity of event data and the limited availability of training datasets [19].
 II-C Image-Event Fusion
-The complementary characteristics of event and image data have been leveraged to enhance various visual tasks, including semantic segmentation [zhang2023cmx, zhang2023cmx], object detection [zhang2021object, zhou2023rgb], and depth estimation [ramnet, srfnet].
-RAMNet [ramnet] employs an RNN-based approach to effectively utilize asynchronous event data, while EVEN [shi2023even] improves low-light performance by fusing modalities prior to inputting them into an MDE network.
-SRFNet [srfnet] guides modality interaction using spatial reliability masks, and PCDepth [pcdepth] integrates features at the modality level for more effective fusion.
+The complementary characteristics of event and image data have been leveraged to enhance various visual tasks, including semantic segmentation [30, 28], object detection [31], and depth estimation [6, 16].
+RAMNet [6] employs an RNN-based approach to effectively utilize asynchronous event data, while EVEN [22] improves low-light performance by fusing modalities prior to inputting them into an MDE network.
+SRFNet [16] guides modality interaction using spatial reliability masks, and PCDepth [13] integrates features at the modality level for more effective fusion.
 Our FUSE is the first to leverage knowledge transfer from an image foundation model for both feature encoding and fusion, significantly enhancing robustness and accuracy in image-event joint depth estimation.
@@
 The high- and low-frequency components are combined through addition and processed by LayerNorm.
-TABLE I: Quantitative results on the MVSEC datast [mvsec].
+TABLE I: Quantitative results on the MVSEC datast [33].
 $\downarrow$ indicates lower is better and denotes higher is better.
@@
 III-B Event Representation
-The event stream is first transformed into a voxel grid representation [ramnet] to make it compatible with existing network architectures.
+The event stream is first transformed into a voxel grid representation [6] to make it compatible with existing network architectures.
 Each event $e_{n}$ is represented by the tuple $(x_{n},y_{n},t_{n},p_{n})$, which means that at time $t_{n}$, the log intensity change at the pixel position $(x_{n},y_{n})$ exceeds a threshold.
@@
 As such, the asynchronous event stream $\mathbf{S}=\{e_{n}\}_{n=0}^{N-1}$ must be transformed into a frame-like representation to be compatible with existing network architectures.
-Following prior works [ramnet], [srfnet], [pcdepth], [e2depth] we convert the temporal stream of events into a voxel grid [gehrig2019end].
+Following prior works [6], [16], [13], [10] we convert the temporal stream of events into a voxel grid [7].
 Given an event stream $\mathit{S}=\{e_{n}\}_{n=0}^{N-1}$ within the time interval $\Delta T=t_{N-1}-t_{0}$, we transform it into a voxel grid with spatial dimensions $H\times W$ and $B$ time bins:
-$\mathit{V}(x,y,t)=\sum p_{i}\delta\left(x-x_{i},y-y_{i}\right)\max\left\{0,1-\left|t-t_{i}^{*}\right|\right\}$ (1)
+$\mathit{V}(x,y,t)=\sum p_{i}\delta\left(x-x_{i},y-y_{i}\right)\max\left\{0,1-\left|t-t_{i}^{*}\right|\right\},$ (1)
 where $t_{i}^{*}=\frac{B-1}{\Delta T}(t_{i}-t_{0})$. The voxel grid $\mathit{V}\in\mathbb{R}^{H\times W\times B}$. We choose $B=3$ time bins.
@@
 Stage I: Parameter-Efficient Feature Alignment: In this stage, we align event data representations with the image domain’s latent space.
-We initialize $\mathcal{E}_{E}$ and $\mathcal{D}_{E}$ using weights from Depth Anything V2 [depthanything2] to inherit its geometric understanding ability.
-To enable efficient adaptation while preventing catastrophic forgetting [forgetcatastrophic], we implement LoRA[lora], where only the LoRA matrices and PatchEmbed Layer parameters (1.8% of $\mathcal{E}_{E}$) are updated.
+We initialize $\mathcal{E}_{E}$ and $\mathcal{D}_{E}$ using weights from Depth Anything V2 [25] to inherit its geometric understanding ability.
+To enable efficient adaptation while preventing catastrophic forgetting [15], we implement LoRA[11], where only the LoRA matrices and PatchEmbed Layer parameters (1.8% of $\mathcal{E}_{E}$) are updated.
 This allows adaptation to event data while retaining the foundation model’s generalization capability.
@@
 III-E) that enforces consistency in both output space (depth prediction) and latent representations.
-Stage II: Robust Feature Fusion: In this stage, we focus on training the FreDFuse (Sec.
-III-D) in the image-event joint estimator to robustly fuse multi-modal features from image and event.
-The image encoder $\mathcal{E}_{I}$ and decoder $\mathcal{D}$ are inherited from the image foundation model, e.g., Depth Anything V2, while the event encoder $\mathcal{E}_{E}$ is obtained from Stage I.
-Only the FreDFuse is trainable during this stage.
-To improve the robustness and generalization of the proposed method, we also generate randomly degraded image-event pairs as inputs during training.
-The degraded inputs incorporate diverse noise types, including brightness adjustment, overexposure, motion blur, and occlusion masking.
-This systematic degradation strategy compels the feature fusion module to identify robust cross-modal relationships that withstand various sensor-specific degradations, thereby enhancing the model’s generalizability and resilience.
-Here, we use the same objective as Eq.
-10 to maintain cross-modal consistency.
+Stage II: Robust Feature Fusion: This stage focuses on training only the FreDFuse module (Sec.
+III-D) for robust multi-modal fusion, while keeping other components frozen: the image encoder ($\mathcal{E}_{I}$) and decoder ($\mathcal{D}$) from a foundation model (e.g., Depth Anything V2), and the event encoder ($\mathcal{E}_{E}$) from Stage I.
+To improve robustness and generalization, we employ a systematic data augmentation strategy by training on randomly degraded image-event pairs.
+This includes global brightness adjustments (factor $\in[0.5,1.5]$) and localized degradations applied to 50% of the pairs in random $20\%\times 20\%$ regions (detailed in Table II).
+This strategy forces the fusion module to learn resilient cross-modal correlations, using the objective from Eq.
+10 to enforce consistency.
+TABLE II: Parameterization of the random degradation strategy.
+  Degradation Type Parameters Applicable Modality Gaussian Blur Kernel size: $25\times 25$, $\sigma=2$ Image & Event Overexposure $\alpha=2.5$, $\beta=200$ Image only Occlusion Fill region with black pixels Image & Event
 III-D Frequency-Decoupled Fusion Module
@@
 ${L}_{\text{align}}={L}_{1}+{L}_{\cos},$ (10)
-where ${L}_{1}=||{d}-{d}^{*}||_{1}$ aligns output predictions, and ${L}_{\cos}=\left[1-\cos(\mathbf{F},\mathbf{F}^{*})\right]\cdot\mathbb{I}({\alpha\leq\cos(\mathbf{F},\mathbf{F}^{*})\leq\beta})$ aligns latent representations.
-Here, $\alpha=0.2$, $\beta=0.85$, and $\mathbb{I}(\cdot)$ ensure loss activation within the specified similarity range, preventing over-regularization.
-Optimization Objective of Target Datasets: When transferring to target datasets with metric depth labels, we replace pseudo-supervision with the Scale-invariant Logarithmic (SiLog) loss [eigen].
+[…]
+Optimization Objective of Target Datasets: When transferring to target datasets with metric depth labels, we replace pseudo-supervision with the Scale-invariant Logarithmic (SiLog) loss [3].
 Given a predicted depth map $\mathbf{d^{*}}$ and depth ground truth $\mathbf{d}$, the loss is computed over the pixels as follows:
@@
 Figure 4: Qualitative analysis of the MVSEC dataset, outdoor_day1 scene.
-(a) and (b) show the input image and event data; (c) depicts the image-event joint estimation by HMNet[hmnet]; (d) shows the event-only estimation by HMNet; (e) presents our proposed FUSE; (f) shows the depth ground truth.
+(a) and (b) show the input image and event data; (c) depicts the image-event joint estimation by HMNet[9]; (d) shows the event-only estimation by HMNet; (e) presents our proposed FUSE; (f) shows the depth ground truth.
 Figure 5: Qualitative analysis of the MVSEC dataset, outdoor_night1 scene.
-(a) and (b) show the input image and event data; (c) depicts the image-event joint estimation by HMNet[hmnet]; (d) shows the event-only estimation by HMNet; (e) presents our proposed FUSE; (f) shows the depth ground truth.
-Figure 6: Qualitative comparison of zero-shot predictions when the image modality is severely degraded.(a) denotes the original image, (b) represents the image degraded by low light, overexposure, and occlusion masks, (c) denotes the event data, and (d) and (e) show the predicted results from DepthAnything V2 [depthanything2] and our FUSE, respectively.
+(a) and (b) show the input image and event data; (c) depicts the image-event joint estimation by HMNet[9]; (d) shows the event-only estimation by HMNet; (e) presents our proposed FUSE; (f) shows the depth ground truth.
+Figure 6: Qualitative comparison of zero-shot predictions when the image modality is severely degraded.(a) denotes the original image, (b) represents the image degraded by low light, overexposure, and occlusion masks, (c) denotes the event data, and (d) and (e) show the predicted results from DepthAnything V2 [25] and our FUSE, respectively.
 IV Experiment
 To validate the effectiveness of our proposed FUSE framework in addressing the scarcity of event-depth data and modality disparity, we conduct comprehensive experiments under strictly controlled conditions.
-Our evaluation protocol emphasizes three critical aspects: (1) Generalization capability across real-world (MVSEC) and synthetic (DENSE) scenarios; (2) Performance constraints with frozen image-event joint encoders; (3) Computational efficiency in multi-modal fusion.
 IV-A Experimental Settings
-We utilize EventScape [ramnet] for image-event pairs in PST (without depth ground truth).
-We evaluate on MVSEC [mvsec] and DENSE [e2depth] datasets, where MVSEC is a real-world scenario while EventScape and DENSE are synthetic scenarios.
-MVSEC Dataset: The MVSEC dataset uses a pair of DAVIS cameras to capture grayscale images and event data at a resolution of $346\times 260$.
-Depth ground truth is recorded using a LiDAR sensor at 20 Hz.
-Grayscale images are captured at 10 Hz during the day and 45 Hz at night.
-Metrics: Following [pcdepth], we evaluate performance using the metrics: Absolute Relative Error (Abs.Rel), Root Mean Square Error (RMSE), Logarithmic Squared Error (RMSELog), average absolute depth errors at different cut-off depth distances (i.e., 10m, 20m, and 30m), and accuracy ($\delta<1.25^{n},n=1,2,3$).
+We utilize EventScape [6] for image-event pairs in PST (without depth ground truth).
+We evaluate on MVSEC [33] and DENSE [10] datasets, where MVSEC is a real-world scenario while EventScape and DENSE are synthetic scenarios.
+Metrics: Following [13], we evaluate performance using the metrics: Absolute Relative Error (Abs.Rel), Root Mean Square Error (RMSE), Logarithmic Squared Error (RMSELog), average absolute depth errors at different cut-off depth distances (i.e., 10m, 20m, and 30m), and accuracy ($\delta<1.25^{n},n=1,2,3$).
 Implementation Details: The model is trained using the Adam optimizer with a learning rate of 5e-5, implemented in PyTorch on two RTX 3090ti GPUs with a batch size of 48.
-PST on EventScape [ramnet] is trained for 10 epochs, and fine-tuning on MVSEC and DENSE for 20 epochs.
+PST on EventScape [6] is trained for 10 epochs, and fine-tuning on MVSEC and DENSE for 20 epochs.
 Manual alignment is needed for the MVSEC dataset due to the asynchronous nature of images, events, and depth labels.
-As in [pcdepth], event data from the first 50 ms prior to each depth ground truth is paired with the most recent image to form the data pairs.
+As in [13], event data from the first 50 ms prior to each depth ground truth is paired with the most recent image to form the data pairs.
 IV-B Evaluation on MVSEC and DENSE Datasets
-We report quantitative results on the real-world MVSEC [mvsec] and synthetic DENSE [e2depth] datasets.
-For MVSEC, we use the outdoor_day2, outdoor_night2, and outdoor_night3 sequences for training, and outdoor_day1 and outdoor_night1 for testing, with depth range from 1.97 to 80 meters [pcdepth].
-For DENSE, we use the provided training and test sets, with depth range from 3.34 to 1000 meters [ramnet, srfnet].
-We freeze the image-event joint encoder and train only the depth decoder.
+We evaluate our method on the real-world MVSEC [33] and synthetic DENSE [10] datasets.
+For MVSEC, we train on outdoor_day2, outdoor_night2, and outdoor_night3, and test on outdoor_day1 and outdoor_night1 (depth range: 1.97-80m) [13].
+For DENSE, we use its official splits (depth range: 3.34-1000m) [6, 16].
+During training, we freeze the joint encoder and only train the depth decoder.
 Analysis on MVSEC Dataset: Tab.
 I presents a quantitative comparison between our method and SOTA methods for event-based estimation and image-event joint estimation on the MVSEC dataset.
-In the outdoor_day1 scene, our method outperforms the best image-event joint estimation approach [pcdepth] by $14\%$ and $10.2\%$ in Abs.Rel and RMSELog, respectively.
+In the outdoor_day1 scene, our method outperforms the best image-event joint estimation approach [13] by $14\%$ and $10.2\%$ in Abs.Rel and RMSELog, respectively.
 The progress achieved by freezing the encoder and dealing with the significant disparity between source and target data clearly demonstrates the effectiveness of our proposed FUSE in addressing the scarcity of event depth data and providing robust and accurate depth estimation.
@@
 4 and Fig.
-5 present qualitative comparisons with HMNet[hmnet] for the outdoor_day1 and outdoor_night1 scenes, respectively.
+5 present qualitative comparisons with HMNet[9] for the outdoor_day1 and outdoor_night1 scenes, respectively.
 FUSE provides finer and more stable depth predictions.
@@
 Analysis on DENSE Dataset: Tab.
-II presents the quantitative results on the synthetic DENSE dataset [e2depth].
+III presents the quantitative results on the synthetic DENSE dataset [10].
 Our method achieves improvements of $24.9\%$ in Abs.Rel and $33.4\%$ in RMSELog over previous approaches.
@@
 Our FUSE is still able to provide stable depth estimation even when some regions of the image are completely destroyed.
-TABLE II: Comparison of performances on synthetic DENSE dataset [e2depth]
+TABLE III: Comparisons on synthetic DENSE dataset [10]
 […]
 IV-C Ablation Studies
-TABLE III: Ablation studies for Image-Event fusion models
-  Knowledge Transformer Feature Fusion Module Trainable Paras (M) Baseline-1 None Cross-Attention 49.2 Baseline-2 None FreDFuse 51.6 Baseline-3 One-stage FreDFuse 6+2.7 FUSE PST FreDFuse 1.4+4.7+2.7
-We perform ablation experiments using the ViT-Small model as the backbone on the MVSEC dataset, as outlined in Tab.III.
-In Baseline-1 and Baseline-2, the model is trained from scratch on image-event pairs with depth ground truth.
-Baseline-1 uses two cross-attention mechanisms for feature fusion, while Baseline-2 incorporates the proposed FreDFuse (Sec.
+TABLE IV: Ablation studies for Image-Event fusion models
+  Knowledge Transformer Fusion Module Paras (M) Baseline-1 None Cross-Attention 49.2 Baseline-2 None FreDFuse 51.6 Baseline-3 One-stage FreDFuse 6+2.7 FUSE PST FreDFuse 1.4+4.7+2.7
+We conduct ablation studies on the MVSEC dataset using the ViT-Small backbone (Tab.
+IV).
+Baseline-1 and Baseline-2 are trained from scratch with depth supervision; the former uses cross-attention for fusion, while the latter adopts our proposed FreDFuse (Sec.
 III-D).
-Baseline-3 combines FreDFuse with a one-stage knowledge transfer approach, initializing the event encoder with image-based foundation model weights and training both the event encoder and FreDFuse simultaneously.
+Baseline-3 adds one-stage knowledge transfer by initializing the event encoder with image-based foundation model weights and training it jointly with FreDFuse.
 Tab.
-IV and Tab.
-V present the results of the ablation experiments on the MVSEC dataset, specifically in the outdoor_day1 and outdoor_night1 scenes.
-The results demonstrate that both components of our proposed method, PST and FreDFuse, positively influence model performance.
+V and Tab.
+VI report results on outdoor_day1 and outdoor_night1.
+Both PST and FreDFuse contribute significantly to performance gains.
 Effectiveness of FreDFuse: In the outdoor_day1 scene, Baseline-2 demonstrates an $8\%$ improvement over Baseline-1 in the Abs.
@@
 This highlights the significance of the two-stage training with degraded image-event pairs for enhancing the model’s generalization and robustness.
-V Conclusion
-In this paper, we propose FUSE, a novel framework that addresses two fundamental challenges in image-event joint depth estimation: cross-modal knowledge transfer under extreme data scarcity and physics-driven frequency conflicts during feature fusion.
-By aligning image-event latent spaces through a two-stage adapter tuning strategy, PST successfully transfers geometric priors from image depth foundation models to the joint modality domain without requiring any depth-labeled image-event pairs.
-This proves that event-depth representations can be effectively bootstrapped from image counterparts when guided by physics-aware constraints, opening new possibilities for data-efficient cross-modal learning.
-Unlike conventional fusion heuristics, the proposed FreDFuse explicitly resolves the spectral incompatibility between static images (low-frequency dominance) and dynamic events (high-frequency bias) through Gaussian-Laplacian pyramid decomposition.
-The FUSE framework achieves SOTA performance on both real-world (MVSEC) and synthetic (DENSE) benchmarks while enabling lightweight deployment.
-While FUSE demonstrates strong zero-shot generalization, two limitations warrant further investigation.
-Current fusion treats events as fixed-interval histograms, potentially underutilizing their microsecond temporal resolution.
-The framework assumes synchronized image-event pairs, which may not hold for all hardware setups.
-Future work will explore event-guided dynamic fusion that adaptively weights frequency branches based on motion intensity, and unsupervised cross-sensor calibration to relax synchronization requirements.
-These extensions could further solidify image-event depth estimation as a universal solution for extreme environment perception.
-TABLE IV: Quantitative results of the ablation experiments on the MVSEC[mvsec] outdoor_day1
+TABLE V: Quantitative results of the ablation experiments on the MVSEC [33] outdoor_day1
 […]
@@
 FUSE 0.719 0.229 6.169 0.294 1.035 1.848 2.267
-TABLE V: Quantitative results of the ablation experiments on the MVSEC[mvsec] outdoor_night1
+TABLE VI: Quantitative results of the ablation experiments on the MVSEC [33] outdoor_night1
 […]
@@
 FUSE 0.613 0.267 6.785 0.352 1.305 1.998 2.513
-[…]
+IV-D Computational Overhead Analysis
+We analyze the computational overhead of our FUSE framework in terms of parameters, GFLOPs, and inference latency, using a ViT-L backbone and $266\times 266$ input resolution on a single RTX 3090 Ti GPU.
+We compare three models: (1) an image-only baseline (Depth Anything V2), (2) a dual-stream baseline with standard cross-attention (Baseline-1), and (3) our FUSE with FreDFuse.
+As shown in Table VII, FUSE roughly doubles the parameters and GFLOPs compared to the image-only baseline due to its dual-encoder design, a common trait in multimodal approaches.
+Compared to Attention Fusion, FreDFuse introduces only a small overhead: +2.56% parameters, +0.46% GFLOPs, but a higher latency (+24.26%), mainly due to its pyramid-based decomposition, which adds memory and synchronization overhead despite low FLOP cost.
+V Conclusion
+We present FUSE, a framework for image-event joint depth estimation that tackles two challenges: cross-modal knowledge transfer and spectral mismatch during fusion.
+A two-stage adapter tuning transfers priors from image depth models to the joint domain without labeled image-event pairs.
+The proposed FreDFuse module leverages Gaussian-Laplacian pyramids to align frequency components from the two modalities.
+FUSE achieves state-of-the-art performance on MVSEC and DENSE benchmarks.
+However, FUSE has several limitations.
+First, converting asynchronous events into fixed-interval frames reduces temporal fidelity, missing fine motion details.
+Second, it processes samples independently, ignoring inter-frame dynamics and temporal consistency.
+Finally, its relatively high latency—due to dual-stream processing and multiscale fusion—limits real-time deployment.
+Future work should focus on two main aspects.
+The first is to explore native asynchronous processing and recurrent architectures to better exploit temporal dynamics and improve depth consistency.
+The second is to optimize the fusion module and investigate model compression techniques, which are crucial for reducing latency and enabling real-time deployment.
+TABLE VII: Efficiency Comparisons Between FUSE and Baselines.
+  Model Paras (M) GFLOPs Latency (ms) Image-Only Baseline 335.32 309.61 29.20 Attention Fusion 660.05 549.21 57.90 FUSE 676.92 551.74 71.95
+ACKNOWLEDGMENT
+The research was supported by the National Natural Science Foundation of China (U23B2009, 62471158).